# Temporal Scaling Law for Large Language Models

### v3

@@ Abstract
Yizhe Xiong1,2 Xiansheng Chen3 Xin Ye3 Hui Chen1,2 Zijia Lin3 Haoran Lian4 Zhenpeng Su5Wei Huang6Jianwei Niu4Jungong Han1Guiguang Ding1,2 1Tsinghua University 2Beijing National Research Center for Information Science and Technology (BNRist) 3Kuaishou Technology 4Beihang University 5Institute of Information Engineering, Chinese Academy of Sciences, Beijing, China 6School of Computer Science, Beijing University of Posts and Telecommunications xiongyizhe2001@gmail.com
@@ 1 Introduction
Differences from Existing Scaling Laws. We summarize the differences between our temporal scaling law and prior scaling laws Kaplan et al. (2020); Henighan et al. (2020); Cherti et al. (2023); Aghajanyan et al. (2023) in Tab. 1. Our temporal scaling law primarily focuses on the evolution of test loss during the pretraining process, while prior scaling laws model the relations of the final test loss and the computational budget (generally decided by model size and training data scale). Our temporal scaling law enables a direct selection of training hyperparameters on a target LLM after the model size and training data scale are decided by prior scaling laws. We underscore that our temporal scaling law is a generic pattern that decoder-based LLMs follow during the pretraining process.
@@ 2.2 Experiment Setup for Deriving Temporal Scaling Law
We use experiment results to derive the temporal scaling law and conduct predictions after deriving parameters of the temporal scaling law from an early training period. Please refer to appendix B for more experiment details.
@@
Models. We train two LLaMA-structure Touvron et al. (2023a) generative language models with 9.8M and 58M parameters to illustrate our temporal scaling law. The architectures of both are identical to the 14M and 70M parameter models in Biderman et al. (2023). The differences between model parameters can be ascribed to the different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU). When utilizing the temporal scaling law for predictions and further applications, we adopt the architectures of the 410M and 1.0B parameter models in Biderman et al. (2023), which leads to 468M and 1.2B models, respectively. To further scale up and generalize the validation of temporal scaling law, we also pre-train: (1) a 6.7B model with the same architecture of LLaMA-7B, (2) a 468M GPT-NeoX model, and (3) an 2B MoE model. See appendix F for results on (2) and (3).
[…]
@@ 2.3 Temporal Scaling Law
The original scaling law Kaplan et al. (2020) has validated that the final test loss for training with different model sizes and data sizes follows a power-law. To validate the effectiveness of the power-law for the temporal evolution of test loss, we pretrain 9.8M and 70M models following Sec. 2.2 and fit the power-law function ($\mathcal{L}=(p_{1}x)^{p_{2}}+p_{3}$, $\{p_{i}\}$ are fitting parameters) to the test loss evolution curve (i.e., treating test loss as a whole to fit in this case) using non-linear least squares. As shown in Tab. 2, though somehow effective, applying the power-law is not precise enough to portray the temporal evolution for the test loss (follow-up experiments in Fig. 3 also demonstrate that directly using the power-law in predicting the future test loss results in larger errors). Such results motivate us to find a better method to understand the test loss evolution during LLM pretraining.
@@
since $t_{n+1}$ has a longer context than $t_{n}$. To illustrate this pattern, we plot the results of test loss for tokens in different positions, as validated on our ID-Val, using both our 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens. We show the results for 400B tokens and leave the rest in the Appendix. As shown in  Fig. 1, both models, although varying in scales, follow the same pattern that test loss for tokens with shorter context is commonly higher than that for tokens with longer context. For both models, the loss values for tokens near the end of the sequence seem to converge at a certain value. We find that this trend can be well fit with a hyperbolic relation:
@@
where $\mathcal{L}_{i}$ is the loss on token position $i(1\leq i\leq n)$, and $a_{0}$, $a_{1}$, and $a_{2}$ are fitting parameters solved by non-linear least squares. The fit for data from our 9.8M and 58M model are presented in Fig. 1. By applying this fit to test loss results of all checkpoints across the training phase, we find that over 99% of them can be fit with such hyperbolic relation with $R^{2}>0.95$, well demonstrating its generality. We term this phenomenon as the dynamic hyperbolic-law. See appendix C for more insights.
We also tested Kaplan’s power-law Kaplan et al. (2020) and other functions that conform to Eq. 4 to fit the curve, but the results are suboptimal. For example, using the power-law yields $R^{2}<0.75$ for over 95% checkpoints. The conclusion still holds on the OOD-Val and on larger-scale models. See detailed results in appendix E.
@@
From the Dynamic Hyperbolic-law to the Overall Temporal Pattern. In the dynamic hyperbolic-law, $a_{2}$ is the converging factor denoting the converged value of loss on each tail token as context length increases. $a_{0}$ denotes the loss gap between the first token and the last token, and $a_{1}$ is the scaling factor along sequence length. To analyze how the derived values of $a_{0}$, $a_{1}$, and $a_{2}$ vary as training steps increase, we collect the derived value tuples of $(a_{0},a_{1},a_{2})$ for all evaluated small model checkpoints during pretraining: $[(a_{0}^{N_{1}},a_{1}^{N_{1}},a_{2}^{N_{1}}),\cdots,(a_{0}^{N_{tot}},a_{1}^{N_{tot}},a_{2}^{N_{tot}})]$, where $N_{i}$ is the number of trained tokens at the $i$-th checkpoint, and $N_{tot}$ is the total number of tokens used for pretraining. Then we plot the collected $(a_{0},a_{1},a_{2})$ values on ID-Val w.r.t. the number of tokens trained (denoted as $N$) for the 58M model in Fig. 2. It can be seen that the evolution of $a_{0}$, $a_{1}$, and $a_{2}$ presents an evident pattern that can be fit by common functions. Therefore, it is possible for us to firstly fit the temporal evolution of $a_{0}$, $a_{1}$, and $a_{2}$, and then fill them into Eq. 5 to depict the temporal evolution of the overall test loss.
@@
where the $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ fitting parameters are solved by non-linear least squares. Observed from the original data in Fig. 2, we find that the value of $a_{0}$ and $a_{1}$ generally converges after a period of training, but suffer from fluctuations after converging due to the uncertainty of the first prediction position in a sequence, which has no previous context to make prediction. To mitigate the fluctuations, we define a separation point ${N}_{sep}$:
@@
[…]
For parameter $a_{2}$, we observe that ${N}_{sep}$ is also a separation point as it marks the shift of its decreasing pattern. Before ${N}_{sep}$, we apply the same function form of the loss gap factor $a_{0}$ as in Eq. 6. After ${N}_{sep}$, we find that its temporal pattern holds strong correlations with the cosine learning rate scheduler, and in turn apply a cosine relation. Similar to Eq. 6, we search the common functions and choose the best fitting function among them for $a_{2}$:
$a_{2}^{N}=\begin{cases}\gamma_{0}\log(\gamma_{1}\log(N)+\gamma_{2})+\gamma_{3},&(N<{N}_{sep})\\ \gamma_{4}\cos(\gamma_{5}N+\gamma_{6})+\gamma_{7},&(N\geq{N}_{sep})\end{cases}$ (9)
@@
We plot the fit of $a_{0}$, $a_{1}$, and $a_{2}$ in Fig. 2. Our fit captures the primary patterns of parameter evolution and ignores the insignificant fluctuations. The conclusion holds on the OOD-Val, and for larger models and models with other structures. See detailed results in appendices E and F.
@@
As listed in Tab. 2, the fit for all ground-truth test losses during the whole training process achieves $R^{2}>0.99$ across different settings, demonstrating the validity of our temporal scaling law. Following Gao et al. (2021), We also validate the fitting results on the LAMBADA Paperno et al. (2016) and the WikiText Merity et al. (2017) datasets, both achieving $R^{2}>0.98$. See appendix E for detailed results.
@@ 2.4 Test Loss Prediction
To predict the overall loss, we first predict the temporal trajectory of $a_{0}^{N}$, $a_{1}^{N}$, and $a_{2}^{N}$. We could fit for $a_{0}^{N}$ and $a_{1}^{N}$ with the functions in Eq. 6:
@@
where $\{\tilde{\alpha}_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are fit using all collected $\{\mathcal{L}_{i}^{N}\}(N<N_{train})$ with non-linear least squares. Following Eq. 7, we could estimate the separation point as $\tilde{N}_{sep}$. Based on the relationship of $N_{train}$ and $\tilde{N}_{sep}$, the test loss is predicted as:
Situation #1 (In most occasions): $N_{train}\leq\tilde{N}_{sep}$, i.e., in an early training stage where the separate point is not reached yet. For $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$, we apply the fit Eq. 11 for $N\leq\tilde{N}_{sep}$. For $\tilde{a}_{2}^{N}$, we first fit the first segment in Eq. 9:
@@
According to the definition in Eq. 7, we simply stabilize the values for $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$ as $\tilde{a}_{0}^{\tilde{N}_{sep}}$ and $\tilde{a}_{1}^{\tilde{N}_{sep}}$ correspondingly for $N>\tilde{N}_{sep}$, fulfilling the derivative condition in Eq. 13 approximately. For $\tilde{a}_{2}^{N}$ at $N\geq\tilde{N}_{sep}$, to eliminate the number of unknown variables, we directly adopt the discovery from fitting Eq. 9 that $a_{2}$ resembles half of a cosine period, and directly set $\tilde{\gamma}_{5}=\frac{\pi}{N_{tot}}$. We set $\tilde{\gamma}_{6}=-\frac{\pi}{N_{tot}}N_{w}$, where $N_{w}$ is the number of tokens used for warmup training, to ensure that $N=N_{w}$ marks the start point of a cosine period. We in turn solve for the rest parameters $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ in the following function based on Eq. 13:
@@
After acquiring the complete estimations of $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$, we could calculate the test loss prediction ~$\tilde{\mathcal{L}}^{N}$ for each training step with Eq. 10.
@@
Apart from the power-law, we also choose the reciprocal ($\mathcal{L}=\frac{a_{0}}{1+a_{1}N}+a_{2}$) and the logarithmic ($\mathcal{L}=\log(a_{1}+a_{2}N)+a_{3}$) relations as baselines that model the test loss as a whole. We predict the subsequent test loss after deriving parameters of the temporal scaling law from, respectively, 10%, 20%, 30%, and 40% of the training process. Note that here we apply the larger models with 468M, 1.2B, and 6.7B parameters for validation. As shown in LABEL:{fig:fit_exp_result}, on different model scales and proportions of training, while the baselines hardly generate reliable results, our method always yields accurate results with $\text{MSE}<10^{-2}$. Our method yields far better $R^{2}$ as well. For example, when using 10% data for 1.2B model on ID-Val, our method yields $R^{2}=0.87$, while the power-law, reciprocal, and logarithmic baselines yield $R^{2}=-1.26,-3.47,-1.02$, correspondingly. Those prediction results provide strong evidence for the reliability of our temporal scaling law. The conclusion holds on the OOD-Val. See appendix G for detailed results.

### v4

@@ Abstract
Yizhe Xiong1,2  Xiansheng Chen3  Xin Ye3  Hui Chen2thanks: Corresponding Authors.  Zijia Lin3∗  Haoran Lian4 Zhenpeng Su5Wei Huang6Jianwei Niu4Jungong Han7Guiguang Ding1,2∗ 1School of Software, Tsinghua University 2BNRist, Tsinghua University 3Kuaishou Technology 4Beihang University 5University of Chinese Academy of Sciences 6School of Computer Science, BUPT 7 Department of Automation, Tsinghua University {xiongyizhe2001, jichenhui2012}@gmail.com linzijia07@tsinghua.org.cn dinggg@tsinghua.edu.cn
@@ 1 Introduction
Differences from Existing Scaling Laws. We summarize the differences between our temporal scaling law and prior scaling laws Kaplan et al. (2020); Henighan et al. (2020); Cherti et al. (2023); Aghajanyan et al. (2023) in Tab.˜1. Our temporal scaling law primarily focuses on the evolution of test loss during the pretraining process, while prior scaling laws model the relations of the final test loss and the computational budget (generally decided by model size and training data scale). Our temporal scaling law enables a direct selection of training hyperparameters on a target LLM after the model size and training data scale are decided by prior scaling laws. We underscore that our temporal scaling law is a generic pattern that decoder-based LLMs follow during the pretraining process.
@@ 2.2 Experiment Setup for Deriving Temporal Scaling Law
We use experiment results to derive the temporal scaling law and conduct predictions after deriving parameters of the temporal scaling law from an early training period. Please refer to appendix˜B for more experiment details.
@@
Models. We train two LLaMA-structure Touvron et al. (2023a) generative language models with 9.8M and 58M parameters to illustrate our temporal scaling law. The architectures of both are identical to the 14M and 70M parameter models in Biderman et al. (2023). The differences between model parameters can be ascribed to the different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU). When utilizing the temporal scaling law for predictions and further applications, we adopt the architectures of the 410M and 1.0B parameter models in Biderman et al. (2023), which leads to 468M and 1.2B models, respectively. To further scale up and generalize the validation of temporal scaling law, we also pre-train: (1) a 6.7B model with the same architecture of LLaMA-7B, (2) a 468M GPT-NeoX model, and (3) an 2B MoE model. Please refer to appendix˜G in the Appendix for training and fitting results on (2) and (3).
Training. Following LLaMA Touvron et al. (2023a), we apply the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of 3ee-4. Following most open-source LLMs Touvron et al. (2023a); Biderman et al. (2023), we use the cosine learning rate decay schedule Loshchilov and Hutter (2017) for all experiments to guarantee the broad applicability of our work. All models are pretrained with 400B tokens and $1k$ total warmup steps.
@@ 2.3 Temporal Scaling Law
The original scaling law Kaplan et al. (2020) has validated that the final test loss for training with different model sizes and data sizes follows a power-law. To validate the effectiveness of the power-law for the temporal evolution of test loss, we pretrain 9.8M and 70M models following Sec.˜2.2 and fit the power-law function ($\mathcal{L}=(p_{1}x)^{p_{2}}+p_{3}$, $\{p_{i}\}$ are fitting parameters) to the test loss evolution curve (i.e., treating test loss as a whole to fit in this case) using non-linear least squares. As shown in Tab.˜2, though somehow effective, applying the power-law is not precise enough to portray the temporal evolution for the test loss (follow-up experiments in Fig.˜3 also demonstrate that directly using the power-law in predicting the future test loss results in larger errors). Such results motivate us to find a better method to understand the test loss evolution during LLM pretraining.
@@
since $t_{n+1}$ has a longer context than $t_{n}$. To illustrate this pattern, we plot the results of test loss for tokens in different positions, as validated on our ID-Val, using both our 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens. We show the results for 400B tokens and leave the rest in the Appendix. As shown in  Fig.˜1, both models, although varying in scales, follow the same pattern that test loss for tokens with shorter context is commonly higher than that for tokens with longer context. For both models, the loss values for tokens near the end of the sequence seem to converge at a certain value. We find that this trend can be well fit with a hyperbolic relation [^2] Note that this hyperbolic relation is actually a special case of the power-law Kaplan et al. (2020); Team et al. (2024) with an exponent of 1-1. Compared to directly applying the power-law, the hyperbolic relation leads to better fitting results. We compare and elaborate this in appendix C of the Appendix.:
@@
where $\mathcal{L}_{i}$ is the loss on token position $i(1\leq i\leq n)$, and $a_{0}$, $a_{1}$, and $a_{2}$ are fitting parameters solved by non-linear least squares. The fit for data from our 9.8M and 58M model are presented in Fig.˜1. By applying this fit to test loss results of all checkpoints across the training phase, we find that over 99% of them can be fit with such hyperbolic relation with $R^{2}>0.95$, well demonstrating its generality. We term this phenomenon as the dynamic hyperbolic-law. See appendix˜C for more insights.
We also tested Kaplan’s power-law Kaplan et al. (2020) and other functions that conform to Eq.˜4 to fit the curve, but the results are suboptimal. For example, using the power-law yields $R^{2}<0.75$ for over 95% checkpoints. The conclusion still holds on the OOD-Val and on larger-scale models. See detailed results in appendix˜F.
@@
From the Dynamic Hyperbolic-law to the Overall Temporal Pattern. In the dynamic hyperbolic-law, $a_{2}$ is the converging factor denoting the converged value of loss on each tail token as context length increases. $a_{0}$ denotes the loss gap between the first token and the last token, and $a_{1}$ is the scaling factor along sequence length. To analyze how the derived values of $a_{0}$, $a_{1}$, and $a_{2}$ vary as training steps increase, we collect the derived value tuples of $(a_{0},a_{1},a_{2})$ for all evaluated small model checkpoints during pretraining: $(a_{0}^{N_{1}},a_{1}^{N_{1}},a_{2}^{N_{1}}),\cdots,(a_{0}^{N_{tot}},a_{1}^{N_{tot}},a_{2}^{N_{tot}})]$, where $N_{i}$ is the number of trained tokens at the $i$-th checkpoint, and $N_{tot}$ is the total number of tokens used for pretraining. Then we plot the collected $(a_{0},a_{1},a_{2})$ values on ID-Val w.r.t. the number of tokens trained (denoted as $N$) for the 58M model in Fig.˜2. It can be seen that the evolution of $a_{0}$, $a_{1}$, and $a_{2}$ presents an evident pattern that can be fit by common functions. Therefore, it is possible for us to firstly fit the temporal evolution of $a_{0}$, $a_{1}$, and $a_{2}$, and then fill them into Eq.˜5 to depict the temporal evolution of the overall test loss.
@@
where the $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ fitting parameters are solved by non-linear least squares. Observed from the original data in Fig.˜2, we find that the value of $a_{0}$ and $a_{1}$ generally converges after a period of training, but suffer from fluctuations after converging due to the uncertainty of the first prediction position in a sequence, which has no previous context to make prediction. To mitigate the fluctuations, we define a separation point ${N}_{sep}$:
@@
Here we empirically set $\epsilon=10^{-4}/\text{1B tokens}$.
For parameter $a_{2}$, we observe that ${N}_{sep}$ is also a separation point as it marks the shift of its decreasing pattern. Before ${N}_{sep}$, we apply the same function form of the loss gap factor $a_{0}$ as in Eq.˜6. After ${N}_{sep}$, we find that its temporal pattern holds strong correlations with the cosine learning rate scheduler, and in turn apply a cosine relation. Similar to Eq.˜6, we search the common functions and choose the best fitting function among them for $a_{2}$:
$a_{2}^{N}=\begin{cases}\gamma_{0}\log(\gamma_{1}\log(N)&\!\!\!\!\!\!+\gamma_{2})+\gamma_{3},\\ &\!\!\!\!\!\!\!(N<{N}_{sep})\\ \gamma_{4}\cos(\gamma_{5}N+\gamma_{6}&\!\!\!\!\!\!)+\gamma_{7},\\ &\!\!\!\!\!\!\!(N\geq{N}_{sep})\end{cases}$ (9)
@@
We plot the fit of $a_{0}$, $a_{1}$, and $a_{2}$ in Fig.˜2. Our fit captures the primary patterns of parameter evolution and ignores the insignificant fluctuations. The conclusion holds on the OOD-Val, and for larger models and models with other structures. See detailed results in appendices˜F and G.
Model Scale LAMBADA Wikitext
9.8M 0.9887 0.9809
58M 0.9805 0.9899
468M 0.9831 0.9920
1.2B 0.9859 0.9930
6.7B 0.9872 0.9927
@@
As listed in Tab.˜2, the fit for all ground-truth test losses during the whole training process achieves $R^{2}>0.99$ across different settings, demonstrating the validity of our temporal scaling law. Following Gao et al. (2021), We also validate the fitting results on the LAMBADA Paperno et al. (2016) and the WikiText Merity et al. (2017) datasets. Specifically, we use the validation and the test splits for the LAMBADA dataset, and use the wikitext-2-v1 domain for the WikiText dataset. As shown in Tab.˜3, all results achieve $R^{2}>0.98$, well demonstrating the effectiveness of our temporal scaling law. Please refer to appendix˜F for more fitting results.
@@ 2.4 Test Loss Prediction
Figure 3: MSE results for predicting the subsequent test loss via the temporal scaling law after completing different proportions of the training process. For the 468M, 1.2B, and 6.7B models, $N_{sep}$ lies approximately at 40% of the entire training schedule, and all presented results require predicting through both fitting stages (i.e., $N<N_{sep}$ and $N\geq N_{sep}$). Note that the y-axis representing the MSE error is in log scale.
To predict the overall loss, we first predict the temporal trajectory of $a_{0}^{N}$, $a_{1}^{N}$, and $a_{2}^{N}$. We could fit for $a_{0}^{N}$ and $a_{1}^{N}$ with the functions in Eq.˜6:
@@
where $\{\tilde{\alpha}_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are fit using all collected $\{\mathcal{L}_{i}^{N}\}(N<N_{train})$ with non-linear least squares. Following Eq.˜7, we could estimate the separation point as $\tilde{N}_{sep}$. Based on the relationship of $N_{train}$ and $\tilde{N}_{sep}$, the test loss is predicted as:
Situation #1 (In most occasions): $N_{train}\leq\tilde{N}_{sep}$, i.e., in an early training stage where the separate point is not reached yet. For $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$, we apply the fit Eq.˜11 for $N\leq\tilde{N}_{sep}$. For $\tilde{a}_{2}^{N}$, we first fit the first segment in Eq.˜9:
@@
According to the definition in Eq.˜7, we simply stabilize the values for $\tilde{a}_{0}^{N}$ and $\tilde{a}_{1}^{N}$ as $\tilde{a}_{0}^{\tilde{N}_{sep}}$ and $\tilde{a}_{1}^{\tilde{N}_{sep}}$ correspondingly for $N>\tilde{N}_{sep}$, fulfilling the derivative condition in Eq.˜13 approximately. For $\tilde{a}_{2}^{N}$ at $N\geq\tilde{N}_{sep}$, to eliminate the number of unknown variables, we directly adopt the discovery from fitting Eq.˜9 that $a_{2}$ resembles half of a cosine period, and directly set $\tilde{\gamma}_{5}=\frac{\pi}{N_{tot}}$. We set $\tilde{\gamma}_{6}=-\frac{\pi}{N_{tot}}N_{w}$, where $N_{w}$ is the number of tokens used for warmup training, to ensure that $N=N_{w}$ marks the start point of a cosine period. We in turn solve for the rest parameters $\tilde{\gamma}_{4}$ and $\tilde{\gamma}_{7}$ in the following function based on Eq.˜13:
@@
After acquiring the complete estimations of $\tilde{a}_{0}^{N}$, $\tilde{a}_{1}^{N}$, and $\tilde{a}_{2}^{N}$, we could calculate the test loss prediction $\tilde{\mathcal{L}}^{N}$ for each training step with Eq.˜10.
@@
Apart from the power-law, we also choose the reciprocal ($\mathcal{L}=\frac{a_{0}}{1+a_{1}N}+a_{2}$) and the logarithmic ($\mathcal{L}=\log(a_{1}+a_{2}N)+a_{3}$) relations as baselines that model the test loss as a whole. We predict the subsequent test loss after deriving parameters of the temporal scaling law from, respectively, 10%, 20%, 30%, and 40% of the training process. Note that here we apply the larger models with 468M, 1.2B, and 6.7B parameters for validation. As shown in LABEL:{fig:fit_exp_result}, on different model scales and proportions of training, while the baselines hardly generate reliable results, our method always yields accurate results with $\text{MSE}<10^{-2}$. Our method yields far better $R^{2}$ as well. For example, when using 10% data for 1.2B model on ID-Val, our method yields $R^{2}=0.87$, while the power-law, reciprocal, and logarithmic baselines yield $R^{2}=-1.26,-3.47,-1.02$, correspondingly. Those prediction results provide strong evidence for the reliability of our temporal scaling law. The conclusion holds on the OOD-Val. See appendix˜I for detailed results.